OpenStudy (anonymous):

An oil company distributes oil in a metal can shaped like a cylinder that has an actual radius of 5.1 cm and a height of 15.1cm. A worker incorrectly measured the radius as 5 cm and the height as 15cm. determine the relative error in calculating the surface area, to the nearest thousandth.

OpenStudy (asnaseer):

do you know how to calculate the surface area of a cylinder?

OpenStudy (anonymous):

sa= 2pir^2+ 2pi r h

OpenStudy (anonymous):

\[2 pir^2+2pirh\]

OpenStudy (asnaseer):

ok, so now introduce a small margin of error in 'r' and 'h' - since both have the same margin of error, lets call it 'e'. then we get:\[sa=2\pi r(r+h)\]and\[sa_e=2\pi (r+e)(r+e+h+e)\]

OpenStudy (asnaseer):

if you expand this and subtract to get \(sa-sa_e\) that will give you your margin of error in terms of 'e'

OpenStudy (anonymous):

im so confused

OpenStudy (asnaseer):

ok, which part confuses you?

OpenStudy (anonymous):

how to find relative error

OpenStudy (asnaseer):

relative error is defined as: (error)/(original value)

OpenStudy (asnaseer):

using the symbols above, it would be:\[\frac{sa-sa_e}{sa}\]

OpenStudy (anonymous):

i have a formula measured - actual/ actual

OpenStudy (asnaseer):

so you could just calculate the surface area \(sa\) using r=5.1 and h=15.1. then calculate the erroneous surface area \(sa_e\) using r=5 and h=15. and then just plug these values into the formulae above.

OpenStudy (asnaseer):

yes - sorry - the relative error is as you say:\[\frac{sa_e-sa}{sa}\]

OpenStudy (asnaseer):

I got the subtraction the wrong way round when I entered it above.

OpenStudy (asnaseer):

if you use the formulae I showed above, the relative error comes out to be:\[\frac{2\pi(r+e)(r+h+2e)-2\pi r(r+h)}{2\pi r(r+h)}=\frac{(r+e)(r+h+2e)-r(r+h)}{r(r+h)}\] r=5.1, h=15.1, e=-0.1

OpenStudy (asnaseer):

this simplifies to:\[\frac{3er+eh+2e^2}{r(r+h)}\]don't worry if you don't understand this way of doing it. you can just calculate the surface areas using the actual and measured values of r and h and then just use the formulae you had - that will also work.

OpenStudy (anonymous):

can you do the work with my formula please

OpenStudy (asnaseer):

why don't you show me what you have calculated?

OpenStudy (anonymous):

m-a\a =(15-5)-(15.1-5.1)\ 15.1-5.1

OpenStudy (asnaseer):

no - that is wrong a = the actual surface area = \(2\pi r(r+h)\) with r=5.1, h=15.1 m = the measured surface area = \(2\pi r(r+h)\) with r=5, h=15

OpenStudy (anonymous):

again confused

OpenStudy (asnaseer):

so first calculate the actual surface area 'a' - what do you get?

OpenStudy (asnaseer):

you can use wolfram alpha to help you with these types of calculations. see here: http://www2.wolframalpha.com/input/?i=2*pi*r%28r%2Bh%29%2C+r%3D5.1%2C+h%3D15.1

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

no

OpenStudy (asnaseer):

ok, first work out the surface area of the can using the actual dimensions for r and h. do you know how to do that?

OpenStudy (anonymous):

nooo

OpenStudy (asnaseer):

you gave the correct formule for surface area above (\(sa=2\pi r^2+2\pi rh\)). are you saying you do not know how to use this formulae to work out the surface area?

OpenStudy (anonymous):

i didnt know i had to get surface area

OpenStudy (asnaseer):

the question you posted states "determine the relative error in calculating the surface area, to the nearest thousandth" so you need to calculate the surface area.

OpenStudy (anonymous):

i dont know how

OpenStudy (asnaseer):

you know the formulae for working it out - correct?

OpenStudy (anonymous):

i dont know how todo the entire thing

OpenStudy (asnaseer):

that is why I am trying to break it down into small steps for you. so first we need to calculate the surface area of the actual can using the formulae that you correctly gave as: sa= 2pir^2+ 2pi r h

OpenStudy (asnaseer):

on the original can, r=5.1, h=15.1, so we get: sa = 2 * PI * (5.1)^2 + 2 * PI * 5.1 * 15.1 you can calculate that number can't you?

OpenStudy (anonymous):

647.29

OpenStudy (asnaseer):

they want the answer to the nearest thousandth, so you will need more decimal places, but the basic answer is correct. ok, next step is to calculate the measured surface area by using r=5, h=15. what do you get for that?

OpenStudy (anonymous):

647.29375

OpenStudy (anonymous):

647.293

OpenStudy (asnaseer):

yes - that is correct for the actual surface area. now what do you get for the measured surface area?

OpenStudy (asnaseer):

I would use 647.29375 until you get to the end to avoid rounding errors

OpenStudy (anonymous):

what do i write on my paper to gt the right anwser

OpenStudy (anonymous):

& idk how to do that

OpenStudy (asnaseer):

first you need to calculate the measured surface area using r=5, h=15 in formulae. i.e.: measured sa = 2 * PI * (5)^2 + 2 * PI * 5 * 15 what answer do you get for this calculation?

OpenStudy (anonymous):

628.31853

OpenStudy (asnaseer):

correct, so now, finally, we can work out the relative error as: relative error = (measured - actual)/actual = (628.31853-647.29375)/647.29375 what do you get for this?

OpenStudy (anonymous):

627.318

OpenStudy (asnaseer):

that is wrong - try again

OpenStudy (asnaseer):

(628.31853-647.29375)/647.29375

OpenStudy (anonymous):

-.029

OpenStudy (asnaseer):

thats it! well done! we finally got there - thanks for trying...

OpenStudy (asnaseer):

a lot of people just up

OpenStudy (asnaseer):

**just give up

OpenStudy (anonymous):

thats the final anwser?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

lol it only took an hour

OpenStudy (asnaseer):

I hope you understood how you got there

OpenStudy (anonymous):

kind of you know i still have a bunch of questions in my packet

OpenStudy (asnaseer):

np - I'll try and help if I can - I need to go and eat something now.

OpenStudy (asnaseer):

please type each question separately on the list to the left.

OpenStudy (asnaseer):

that way you will get more people helping you

OpenStudy (anonymous):

k thanks bye

OpenStudy (asnaseer):

bye